OpenStudy (anonymous):

hello, i dont know how to calculate the pd , can you help me to solve this question:- What is the pd across an immersion heater element which changes 3.6 x 103J of electrical energy into heat every second and takes a current of 15 A?

OpenStudy (anonymous):

pd is the potential difference (voltage) across the element

OpenStudy (anonymous):

\[Power =IV =\frac{ QV }{ t } = \frac{ Workdone }{ time }\]

OpenStudy (anonymous):

Workdone on heating element = amount of electrical energy turned into heat

OpenStudy (anonymous):

Now we can use that equation to find the p.d across the element each second \[P = \frac{ 3.6 \times 10^3 J }{ 1 s } = 15A \times V \rightarrow V = 240V\]